OpenStudy (anonymous):

Find an equation of the line containing the given pair of points. (-7,0) and (0, 9). What is the equation of the line? y=____.

hero (hero):

Do you know the slope formula to find the slope, m, given two points?

OpenStudy (anonymous):

hang on, let me go see..

OpenStudy (anonymous):

it doesn't show, this is the question, isn't the (-7, 0) and (0, 9) are the given points? confused!!

OpenStudy (anonymous):

it also just says to type answer in a slope-intercept form. ause integers or fractions for any numbers in the expression.

OpenStudy (anonymous):

*use

hero (hero):

Okay, to find the slope between two points you use the formula: \[m = \frac{y_2 - y_1}{x_2 - x_1}\] where \[x_1 = -7, x_2 = 0, y_1 = 0, y_2 = 9\]

OpenStudy (anonymous):

y=9?

hero (hero):

Once you find the slope, use the slope intercept formula: \[y - y_1 = m(x - x_1)\] Input the coordinates first given point along with the slope you found using the slope formula into the slope intercept formula. Then transform the formula into the form y = mx+b

hero (hero):

You want to find the slope, m, first

OpenStudy (anonymous):

y=\[9/7x+9\]

OpenStudy (anonymous):

sorry, I'm confused...not even sure if i got that right, but i come upn with y=9

hero (hero):

\[y = \frac{9x}{7} + 9\], so your equation is correct. I don't know why you keep saying y = 9

OpenStudy (anonymous):

oh ok, so that would be my answer..not the y=9..i don't know why i keep saying that..sorry. just confused on some of these problems..thanks for you help Hero.

hero (hero):

yw